# Output Prediction of Quantum Circuits based on Graph Neural Networks

### v2

@@ 1 Introduction
Quantum computing, as a new computational paradigm, has the potential to address problems beyond the capabilities of classical computing with greater efficiency and higher speed. It has demonstrated exponential or polynomial advantages in various domains, including combinatorial optimization [1, 2, 3], large-scale communication systems [4, 5, 6, 7], molecular dynamics [8, 9], quantum chemistry[10, 11, 12], and machine learning [13, 14, 15, 16, 17, 18, 19]. Driven by breakthroughs in physical implementation technologies, quantum hardware has undergone rapid development over the past two decades, with numerous quantum computing systems now available [20, 21, 22, 23]. Despite its promising prospects, quantum computing must navigate an extended era of Noisy Intermediate-Scale Quantum (NISQ) devices before entering the fault-tolerant quantum computing era [24, 25]. During this stage, the limitations imposed by coherence times and quantum gate errors pose significant bottlenecks to achieving quantum advantage. Quantum devices currently encounter limitations in resource accessibility, such as runtime and qubit count, which do not match the convenience provided by classical computers. Additionally, classical simulation methods often suffer from inefficiencies. Consequently, some advancements must proceed without adequate benchmark data, significantly impeding the engineering of quantum devices. Against this backdrop, accurately predicting the output of quantum circuits is not only a theoretical challenge but also a crucial step toward optimizing quantum algorithms and enhancing hardware performance. However, due to the complex noise characteristics of quantum circuits and the computational challenges that grow with circuit size, existing classical simulation methods often struggle to provide efficient solutions.
@@
Notably, as it can provide low-cost predictions of specific aspects of quantum circuit outputs once the model is trained, machine learning has recently offered an attractive alternative to direct classical simulation. For instance, in [26], a meticulously designed qubit-scalable Convolutional Neural Network (CNN) is employed to predict expectation values of circuits, delivering results that, under specific conditions, outperform those obtained from the freely available Noisy Intermediate-Scale Quantum (NISQ) devices. In [27], neural networks based on simple Multilayer Perceptrons (MLPs) are used to predict kernel-target alignment (KTA), a cost-effective proxy for quantum kernel classification accuracy, facilitating the identification of high-accuracy quantum kernel circuits. Despite these advancements, CNN-based and MLP-based approaches face limitations in capturing the structural information of quantum circuits, particularly for large-scale systems or circuits with complex topologies. Graph-based models, on the other hand, naturally leverage the directed acyclic graph (DAG) representation of quantum circuits, making them a promising alternative for output prediction. For instance, in [28], Graph Transformers are utilized to predict the Probability of Successful Trials (PST), which positively correlates with circuit fidelity, achieving remarkable performance. Beyond these applications, estimating or predicting quantum circuit outputs still holds broad prospects. For example, circuit output predictors could be integrated into diffusion models, such as those discussed in [29], to guide the generation of high-performance circuits. Alternatively, these predictors could support reinforcement learning (RL) processes, as outlined in [30], by evaluating reward functions, thereby reducing resource consumption during training.
@@
Based on the considerations outlined above, we aim to design a Graph Neural Networks (GNNs) predictor to predict the specific outputs of quantum circuits. This is not only because GNNs can capture the structural information of quantum circuits but also because they can construct individual feature vectors for each node, allowing the embedding of various properties of quantum gates, such as gate type, gate error, and target qubits. In this work, we primarily predict the single-qubit expectation values, two-qubit expectation values, and the overall property of quantum circuits.
@@
To begin with, we propose a GNN-based framework for predicting single-qubit and two-qubit expectation values of quantum circuits, as illustrated in figure 1. These limited expectation values are sufficient for circuits that require only one or two possible output bitstrings, such as the Bernstein–Vazirani (BV) algorithm [31] and the Deutsch–Jozsa algorithm [32]. Utilizing a parameter-free quantum gate set, we randomly generate quantum circuits under both noiseless and noisy conditions. Classical simulations are conducted using the Qiskit library [33] to obtain single-qubit and two-qubit expectation values for our dataset. During the transformation of random circuits into graphs, we design feature vectors for each node to encapsulate properties such as gate type, target qubits, and noise information. Our experiments demonstrate that the GNNs achieve strong predictive performance on datasets with a relatively small circuit search space. In comparison to convolutional neural networks (CNNs)-based methods, our GNNs-based method exhibits superior performance, greater flexibility in handling input circuit structures, and enhanced scalability.
@@
Furthermore, we extend our GNN-based prediction framework in two significant ways. First, we incorporate parameterized quantum gates, expanding its applicability to parameterized quantum circuits (PQCs). Second, instead of solely predicting expectation values, we shift our focus to evaluating the overall property of quantum circuits, specifically predicting the ground state energy of a 4-qubit hydrogen molecule. To compare the performance of PQCs, we propose two schemes: Direct Comparison, which directly predicts the relative performance difference between circuits, and Indirect Comparison, which involves predicting absolute ground state energies before comparison. Our simulation results demonstrate that the Direct Comparison scheme significantly outperforms the Indirect Comparison scheme, offering an average improvement of 36.2% on the same dataset. These findings highlight the potential of GNNs in efficiently evaluating PQC performance, providing a novel and effective perspective for analyzing quantum circuits.
@@ 3.1 GNNs for expectation values prediction
Due to the limited quantum resources of the NISQ devices, it is a meaningful thing to predict the output of the circuit in advance before submitting it for execution. If the predicted output of the circuit is far below a threshold, running this circuit on a real quantum computation will not produce the desired results. To accurately predict circuit output, especially in the presence of noise, we propose a framework for predicting quantum circuits with GNNs, as shown in figure 1.
This is the data-driven approach, and intuitively, estimating expectation values does not require computing the complete density matrix exactly. Therefore, data-driven approaches present the opportunity to provide sufficiently accurate predictions at a lower computational cost. The framework for predicting expectation values is divided into three primary stages. The first stage is to generate random quantum circuits with the assistance of a universal gate set and measure the circuit expectation values, which are used as training data for the GNNs. The second stage is to transform the generated random circuits into a graph structure and formulate feature vectors for each node that can contain device noise information. In the third stage, the graph is processed to predict circuit expectation values using GNNs.
@@ 3.1.1 Random circuits generation
To realize the first stage in our framework, random quantum circuits are generated to provide a database for subsequent GNNs training. We select two single-qubit gates, namely the $\pi/8$ gate ($T$) and Hadamard gate ($H$), and one two-qubit gate CNOT, as the universal gate set to generate random circuits. The universal gate set consisting of parameter-free gates is denoted $S=\left\{T,H,CNOT\right\}$. In the first step, gates are randomly selected from a universal gate set $S$ to create an initial version of the random circuit with $N$ qubits and $P$ layers of gates, which we define as the circuit depth. The integer $N$ corresponds to the number of qubits, and $P$ denotes the number of gates on each qubit. To ensure circuits integrity, we set the initial state of all qubits to $\left|0\right\rangle$, and the circuits include measurement operations. After completing the circuit construction, an optimization strategy is used to eliminate duplicate gates to obtain circuits for later training. These optimization strategies include the equivalence of two consecutive H-gates to an Identity gate, the equivalence of eight consecutive T-gates to an Identity gate, and the elimination of two consecutive CNOTs, etc. The specific process of generating random circuits is illustrated in figure 2.
@@ 3.1.4 Target values
Where $U$ denotes the unitary operator corresponding to the circuit, and the tensor product $\left|0\right\rangle^{\otimes N}\equiv\left|0\right\rangle\otimes\dots\otimes\left|0\right\rangle$ is the initial state. We define $Z_{i}$ as the Pauli operator $Z$ acting on qubit $i$, $i\in\left\{1,\dots,N\right\}$ The single-qubit expectation values can be expressed as
@@
Where $z_{ij}$, $i,j=1,2,\dots,N$ represents the two-qubit expectation values. Certainly, the description provided by the expectation values of single-qubit and two-qubit Pauli-Z measurements is limited. However, these pieces of information are sufficient for certain specific algorithms, such as the Deutsch–Jozsa algorithm [32]. This motivation underlies our exploration of expectation values prediction.
@@ 3.1.5 Construction of GNNs
By leveraging the attention mechanism, the weights of each edge in the graph are dynamically determined, enabling the neural network to better capture information across the entire circuit. The activation function between layers is implemented using the LeakyReLU function with a coefficient of 0.02. Following the feature extraction by the GNNs, we select features from key nodes (excluding input nodes and measurement nodes) and perform average pooling. Subsequently, the pooled feature vector, along with global information about the circuit, including circuit depth, the total number of gates, and the count of CNOT, is fed into a multi-layer fully connected neural network. Then this network outputs the predicted expectation value.
@@ 3.2 GNNs for circuit performance comparison
$E_{VQE}=\min_{\bm{\theta}}\left\langle\bm{0}\left|U^{\dagger}\left(\bm{\theta}\right)HU\left(\bm{\theta}\right)\right|\bm{0}\right\rangle$ (8)
where $H$ denotes the Hamiltonian operator corresponding to the lowest eigenvalue, and $U\left(\bm{\theta}\right)$ denotes the parameterized ansatz, which also represents the PQC in the process of circuit design and implementation. By using the ground state energy values calculated through the Variational Quantum Eigensolver (VQE) as performance metrics for different circuits, and employing these energy values as data to train GNNs, we can predict the performance of randomly generated circuits. This approach lays the foundation for selecting high-performance circuits. In the following two subsections, we will present this work. Section 3.2.1 utilizes the GNNs to predict the ground state energy of a molecule initially and compares the performance of PQCs based on the predicted ground state energy values. Section 3.2.2 builds upon the foundation of Section 3.2.1 by introducing an innovative approach: two different PQCs are simultaneously input into the GNN, which directly predicts the relative probability of their performance comparison.
@@ 3.2.1 Indirectly comparing circuit performance
Differing from the method of generating random circuits in Section 3.1.1, we use $S^{\ast}=\left\{R_{x}\left(\theta\right),R_{y}\left(\beta\right),CNOT\right\}$ as the universal gate set. We randomly generate PQCs with three layers, each layer consisting of a $R_{x}$ layer, a CNOT layer, and a $R_{y}$ layer, as shown in figure 5.
Figure 5: Randomly generated three-layer parameterized quantum circuit (PQC). Each layer has a certain probability of excluding some quantum gates, with the quantum gates enclosed in dashed boxes representing those not selected.
@@
During the training process, the noise model used is derived from IBM’s quantum device, IBM Lagos. When performing pooling in the GNNs, key nodes are selected from each circuit’s corresponding nodes for average pooling, and the results, along with global information including the differences in the number of quantum gates, CNOT gates, and circuit depths between the two circuits, are then fed into the multi-layer fully connected neural networks. Furthermore, the output of the GNNs is the probability that the first circuit outperforms the second circuit, ranging from 0 to 1.

### v3

@@ 1 Introduction
Quantum computing, as a new computational paradigm, has the potential to address problems beyond the capabilities of classical computing with greater efficiency and higher speed. It has demonstrated exponential or polynomial advantages in various domains, including combinatorial optimization [1, 2, 3], large-scale communication systems [4, 5, 6, 7], molecular dynamics [8, 9], quantum chemistry[10, 11, 12], and machine learning [13, 14, 15, 16, 17, 18, 19]. Driven by breakthroughs in physical implementation technologies, quantum hardware has undergone rapid development over the past two decades, with numerous quantum computing systems now available [20, 21, 22, 23]. Despite its promising prospects, quantum computing must navigate an extended era of Noisy Intermediate-Scale Quantum (NISQ) devices before entering the fault-tolerant quantum computing era [24, 25, 26]. During this stage, the limitations imposed by coherence times and quantum gate errors pose significant bottlenecks to achieving quantum advantage. Quantum devices currently encounter limitations in resource accessibility, such as runtime and qubit count, which do not match the convenience provided by classical computers. Additionally, classical simulation methods often suffer from inefficiencies. Consequently, some advancements must proceed without adequate benchmark data, significantly impeding the engineering of quantum devices. Against this backdrop, accurately predicting the output of quantum circuits is not only a theoretical challenge but also a crucial step toward optimizing quantum algorithms and enhancing hardware performance. However, due to the complex noise characteristics of quantum circuits and the computational challenges that grow with circuit size, existing classical simulation methods often struggle to provide efficient solutions.
@@
Notably, as it can provide low-cost predictions of specific aspects of quantum circuit outputs once the model is trained, machine learning has recently offered an attractive alternative to direct classical simulation. For instance, in [27], a meticulously designed qubit-scalable Convolutional Neural Network (CNN) is employed to predict expectation values of circuits, delivering results that, under specific conditions, outperform those obtained from the freely available Noisy Intermediate-Scale Quantum (NISQ) devices. In [28], neural networks based on simple Multilayer Perceptrons (MLPs) are used to predict kernel-target alignment (KTA), a cost-effective proxy for quantum kernel classification accuracy, facilitating the identification of high-accuracy quantum kernel circuits. Despite these advancements, CNN-based and MLP-based approaches face limitations in capturing the structural information of quantum circuits, particularly for large-scale systems or circuits with complex topologies. Graph-based models, on the other hand, naturally leverage the directed acyclic graph (DAG) representation of quantum circuits, making them a promising alternative for output prediction. For instance, in [29], Graph Transformers are utilized to predict the Probability of Successful Trials (PST), which positively correlates with circuit fidelity, achieving remarkable performance. More recently, [30] proposed a general framework of machine learning for quantum error mitigation (ML-QEM), which employs various machine learning models including Linear Regression, Random Forest (RF), and Graph Neural Networks (GNNs) to regress the noisy expectation values obtained from quantum processors toward their approximated noiseless counterparts for error mitigation. Compared with traditional mitigation methods, ML-QEM can significantly reduce the overall cost of quantum error mitigation. Beyond these applications, estimating or predicting quantum circuit outputs still holds broad prospects. For example, circuit output predictors could be integrated into diffusion models, such as those discussed in [31], to guide the generation of high-performance circuits. Alternatively, these predictors could support Reinforcement Learning (RL) processes, as outlined in [32], by evaluating reward functions, thereby reducing resource consumption during training.
@@
Based on the considerations outlined above, we aim to design a Graph Neural Networks (GNNs) predictor to predict the specific outputs of quantum circuits. Within this approach, the GNNs act as a surrogate model that enables large-scale and rapid performance prediction once trained, eliminating the need for repeated costly simulations or hardware executions. In addition to computational efficiency, GNNs also contribute to improving prediction accuracy, as they can effectively capture the intrinsic structural dependencies within quantum circuits. This advantage arises not only from their ability to represent circuit topology more faithfully but also from their capability to construct individual feature vectors for each node, allowing the embedding of various properties of quantum gates, such as gate type, gate error, and target qubits. In this work, we primarily predict the single-qubit expectation values, two-qubit expectation values, and the overall property of quantum circuits.
@@
To begin with, we propose a GNN-based framework for predicting expectation values of quantum circuits, as illustrated in figure 1. These limited expectation values are sufficient for circuits that require only one or two possible output bitstrings, such as the Bernstein–Vazirani (BV) algorithm [33] and the Deutsch–Jozsa algorithm [34]. Utilizing a parameter-free quantum gate set, we randomly generate quantum circuits under both noiseless and noisy conditions. Classical simulations are conducted using the Qiskit library [35] to obtain single-qubit and two-qubit expectation values for our dataset. During the transformation of random circuits into graphs, we design feature vectors for each node to encapsulate properties such as gate type, target qubits, and noise information. Our experiments demonstrate that the GNNs achieve strong predictive performance on datasets with a relatively small circuit search space. For a fair comparison with the well-designed and representative convolutional neural networks (CNNs)-based method [27], which possesses strong structured feature extraction capability, our GNNs-based method demonstrates superior performance, greater flexibility in handling input circuit structures, and improved scalability.
@@
Furthermore, we extend our GNN-based prediction framework in two significant ways. First, we incorporate parameterized quantum gates, expanding its applicability to parameterized quantum circuits (PQCs). Second, instead of solely predicting expectation values, we shift our focus to evaluating the overall property of quantum circuits, specifically predicting the ground-state energy of the H2 molecule encoded on 4 qubits using a standard VQE workflow. In parallel, recent studies have employed deep neural networks to predict VQE circuit parameters directly from molecular geometries, thereby bypassing expensive variational optimization [36, 37]. Such ML-assisted VQE schemes have demonstrated accurate ground-state energy predictions, including for systems beyond H2 (e.g., LiH and BeH2), reinforcing the relevance of ML for PQC-based quantum chemistry. To compare the performance of PQCs, we propose two schemes: Direct Comparison, which directly predicts the relative performance difference between circuits, and Indirect Comparison, which involves predicting absolute ground state energies before comparison. Our simulation results demonstrate that the Direct Comparison scheme significantly outperforms the Indirect Comparison scheme, offering an average improvement of 36.2% on the same dataset. These findings highlight the potential of GNNs in efficiently evaluating PQC performance, providing a novel and effective perspective for analyzing quantum circuits.
@@ 3.1 GNNs for expectation values prediction
Due to the limited quantum resources of the NISQ devices, it is a meaningful thing to predict the output of the circuit in advance before submitting it for execution. If the predicted output of the circuit is far below a threshold, running this circuit on a real quantum computation will not produce the desired results. To accurately predict circuit output, especially in the presence of noise, we propose a framework for predicting quantum circuits with GNNs, as shown in figure 1. This is the data-driven approach, and intuitively, estimating expectation values does not require computing the complete density matrix exactly. Therefore, data-driven approaches present the opportunity to provide sufficiently accurate predictions at a lower computational cost. The framework for predicting expectation values is divided into three primary stages. The first stage is to generate random quantum circuits with the assistance of a universal gate set and measure the circuit expectation values, which are used as training data for the GNNs. The second stage is to transform the generated random circuits into a graph structure and formulate feature vectors for each node that can contain device noise information. In the third stage, the graph is processed to predict circuit expectation values using GNNs.
@@ 3.1.1 Random circuits generation
To realize the first stage in our framework, random quantum circuits are generated to provide a database for subsequent GNNs training. We select two single-qubit gates, namely the $\pi/8$ gate ($T$) and Hadamard gate ($H$), and one two-qubit gate CNOT, as the universal gate set to generate random circuits. The universal gate set consisting of parameter-free gates is denoted $S=\left\{T,H,CNOT\right\}$. In the first step, gates are randomly selected from a universal gate set $S$ to create an initial version of the random circuit with $N$ qubits and $P$ layers of gates, which we define as the circuit depth. The integer $N$ corresponds to the number of qubits, and $P$ denotes the number of gates on each qubit, where two-qubit gates are considered as separate single-qubit operations acting on both qubits. To ensure circuits integrity, we set the initial state of all qubits to $\left|0\right\rangle$, and the circuits include measurement operations. After completing the circuit construction, an optimization strategy is used to eliminate duplicate gates to obtain circuits for later training. These optimization strategies include the equivalence of two consecutive H-gates to an Identity gate, the equivalence of eight consecutive T-gates to an Identity gate, and the elimination of two consecutive CNOTs, etc. The motivation for this step is to ensure that the circuits in our training dataset remain as compact as possible, so that the neural network can focus on learning meaningful structural differences rather than spending capacity on trivial equivalence transformations. The specific process of generating random circuits is illustrated in figure 2. The quantum gates enclosed by the red dashed box in figure 2(a) are the optimizable gates, and the optimized circuit is shown in figure 2(b).
@@ 3.1.4 Target values
Where $U$ denotes the unitary operator corresponding to the circuit, and the tensor product $\left|0\right\rangle^{\otimes N}\equiv\left|0\right\rangle\otimes\dots\otimes\left|0\right\rangle$ is the initial state. We define $Z_{i}$ as the Pauli operator $Z$ acting on qubit $i$, $i\in\left\{1,\dots,N\right\}$ . The single-qubit expectation values can be expressed as
@@
Since $Z_{i}=\left|0\right\rangle_{i}\left\langle 0\right|_{i}-\left|1\right\rangle_{i}\left\langle 1\right|_{i}$, the expectation value of $\left\langle Z_{i}\right\rangle$ can thus be calculated as
$\left\langle Z_{i}\right\rangle=\left|\left\langle\bm{\varphi}|0\right\rangle_{i}\right|^{2}-\left|\left\langle\bm{\varphi}|1\right\rangle_{i}\right|^{2}$ (6)
@@
Where $z_{ij}$, $i,j=1,2,\dots,N$ represents the two-qubit expectation values. Certainly, the description provided by the expectation values of single-qubit and two-qubit Pauli-Z measurements is limited. However, these pieces of information are sufficient for certain specific algorithms, such as the Deutsch–Jozsa algorithm [34]. For instance, in the Deutsch–Jozsa algorithm, the goal is to determine whether a Boolean function $f:\left\{0,1\right\}^{\left(N-1\right)}\longrightarrow\left\{0,1\right\}$ is constant or balanced. The algorithm measures the first N1N-1 qubits, and if the outcome corresponds to the bit string $00\cdots 0$ the function is constant; otherwise, it is balanced. In other words, predicting the expectation values of the N1N-1 single-qubit observables $\left\langle Z_{i}\right\rangle$ (for $i=1,2,\cdots,N-1$) is already sufficient to infer the result, that is, if $\left\langle Z_{i}\right\rangle=1$ for all $i=1,2,\cdots,N-1$ , the output bitstring must be $00\cdots 0$, indicating a constant function. Conversely, the function is balanced. This motivation underlies our exploration of expectation values prediction.
@@ 3.1.5 Construction of GNNs
By leveraging the attention mechanism, the weights of each edge in the graph are dynamically determined, enabling the neural network to better capture information across the entire circuit. After feature extraction by the GNNs, features from key nodes (excluding input and measurement nodes) are selected and subjected to average pooling to obtain a 31-dimensional pooled vector. Meanwhile, the global circuit information, including the total number of gates, the number of CNOT gates, and the circuit depth, is fed into a two-layer fully connected network, where both the hidden and output layers have a dimensionality of 12. These global features are then concatenated with the pooled feature vector. Finally, this concatenated vector is processed through a three-layer fully connected network with hidden layer dimensions of 256, 128, and 64 to produce the final prediction result. The activation function between layers is implemented using the LeakyReLU function with a coefficient of 0.02.
@@ 3.2 GNNs for circuit performance comparison
[…]
where $H$ denotes the Hamiltonian operator corresponding to the lowest eigenvalue, and $U\left(\bm{\theta}\right)$ denotes the parameterized ansatz, which also represents the PQC in the process of circuit design and implementation. By using the ground state energy values calculated through the Variational Quantum Eigensolver (VQE) as performance metrics for different circuits, and employing these energy values as data to train GNNs, we can predict the performance of randomly generated circuits. Specifically, we use the commonly adopted VQE optimizer COBYLA to optimize the circuit parameters. COBYLA is a gradient-free optimization method that does not require explicit computation of parameter gradients. The number of optimization iterations is set to 200, after which the optimized ground-state energy is obtained and used as the training target for the GNNs. It should be noted that the GNNs do not participate in the VQE optimization process. The VQE is solely employed to generate ground-state energy data for different circuits, which are used as labels for training the GNNs. It is worth noting that the predicted energy values are jointly affected by both the circuit structure and the optimization process. Although our model does not explicitly simulate the optimization process itself, it indirectly reflects the quality of circuit structures by learning the final energies of different circuits optimized under a fixed optimizer. This approach lays the foundation for selecting high-performance circuits. In the following two subsections, we will present this work. Section 3.2.1 utilizes the GNNs to predict the ground state energy of a molecule initially and compares the performance of PQCs based on the predicted ground state energy values. Section 3.2.2 builds upon the foundation of Section 3.2.1 by introducing an innovative approach: two different PQCs are simultaneously input into the GNN, which directly predicts the relative probability of their performance comparison.
@@ 3.2.1 Indirectly comparing circuit performance
Differing from the method of generating random circuits in Section 3.1.1, we use $S^{\ast}=\left\{R_{x}\left(\theta\right),R_{y}\left(\beta\right),CNOT\right\}$ as the universal gate set. We randomly generate PQCs with three layers, each layer consisting of a $R_{x}$ layer, a CNOT layer, and a $R_{y}$ layer. A single layer of the parameterized quantum circuit (PQC) is randomly generated as shown in figure 5. To construct a three-layer PQC, the process shown in figure 5 is repeated three times, and the resulting three circuits are concatenated to form a complete quantum circuit.
Figure 5: Randomly generated single layer parameterized quantum circuit (PQC) with the quantum gates enclosed in dashed boxes representing those not selected.
@@
During the graph construction process, two quantum circuits are merged into a single graph consisting of two disconnected subgraphs, as illustrated in the upper part of figure 6. In the four-layer Graph Neural Networks (GNNs), each of the first three layers performs not only node-wise message aggregation but also average pooling on the key nodes (excluding input and measurement nodes) within each subgraph. The pooled feature vectors from the two subgraphs are then subtracted to obtain the difference features. Differently from the previous layers, the final layer also performs average pooling on the key nodes of both subgraphs but directly concatenates the two pooled feature vectors instead of taking their difference, aiming to preserve more individual information from each subgraph after aggregation. The difference vectors obtained from the first three layers and the concatenated feature vector from the final layer are then concatenated to form the output representation of the GNNs. Meanwhile, global information, including the differences in the total number of quantum gates, CNOT gates, and circuit depths between the two circuits, is first passed through two fully connected layers. The resulting output is then concatenated with the GNNs output to form a unified feature vector, which is subsequently fed into three fully connected layers to produce the final prediction. During the training process, the noise model used is derived from IBM’s quantum device, IBM Lagos. Furthermore, the output of the GNNs is the probability that the first circuit outperforms the second circuit, ranging from 0 to 1. To summarize the distinction between the two comparison schemes, in the indirect comparison scheme, the GNNs predict the ground-state energy obtained from VQE optimization for each circuit, whereas in the direct comparison scheme, the inputs are two circuit graphs, and the output represents the probability that the first circuit performs better than the second one. The labels are constructed based on the energy difference between the two circuits after Min–Max normalization.